I am currently learning Systems of Linear Equations and Inequalities the problem I am currently working on is:
3x+5y=14
x-2y=-10
I have no idea where to start, or how to work the problem

Answer by fcabanski(1391)

You can use substitution. Solve the second equation for x:
x=2y-10 | Substitute that x value into the first equation and solve for y:
3(2y-10)+5y=14 |Distribute the 3:
6y-30+5y=14 | Combine the y's, add 30 to both sides:
11y=44 | divide both sides by 11:
y = 4 |substitute that into the second equation that you solved for x:
x=2(4)-10 = 8-10 = -2

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com